# A Clique Partitioning-Based Algorithm for Graph Compression

### v2

@@ 2 Background and Motivation
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$$z_{1}$
Figure 1: (a) Given bipartite graph $G(U,W,E)$; (b) Neighborhood tree of vertex $u_{2}\in U$ that shows the path $\omega$ taken from root to the vertex $w_{j}\in W$ at the leaf and number of edges $d_{u_{2},\omega}$ for each node using the tuple $(\omega,d_{u_{2},\omega})$; and (c) the tripartite graph that replaces the $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in the compressed graph $G^{*}(U,W,Z,E^{*})$.
The FM algorithm iteratively extracts $\delta$-cliques from the input graph $G(U,W,E)$ by selecting $\hat{k}=k(n,m,\delta)$ vertices based on the neighborhood trees associated with each of the vertices $u_{i}\in U$. A neighborhood tree for a vertex $u_{i}\in U$ is a binary tree whose nodes at level $i$ correspond to a partition of $W$ into sets of size $|W|/2^{i}$. Thus, the root node corresponds to the set of all the vertices in $W$, while each leaf node corresponds to a vertex in $W$. Figure 1(b) shows the neighborhood tree for vertex $u_{2}\in U$ which has six neighbour vertices in $W$. The neighborhood trees are described in more detail in Appendix A.1. Once the cliques are extracted, FM compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph $G^{*}$ (Figure 1(c)), where each vertex in $Z$ corresponds to one $\delta$-clique extracted by the algorithm. Each vertex of the left and right partitions of the $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The obtained compressed graph $G^{*}$ preserves the path information of the original graph, i.e., the connectivity between the vertices $u_{i}\in U$ and $w_{j}\in W$ of graph $G$. The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$. Feder and Motwani [9] also showed that their FM algorithm can be extended to the case of non-bipartite graphs to obtain compression. In Appendix A.1 we provide a more detailed description of the FM algorithm.
2.2 Motivation.
While the FM algorithm provides a solid foundation for lossless graph compression, there are several aspects that can be improved to achieve better compression and efficiency. In this section, we motivate our approach by delineating three critical aspects of a $\delta$-clique-based graph compression technique aimed at improving both compression ratio and execution efficiency.
Vertex selection. For each $\delta$-clique $C_{q}$, the FM algorithm iteratively chooses $\hat{k}$ vertices from $W$ to form the right partition $W_{q}$ of $C_{q}$ using the neighborhood trees associated to each $u_{i}\in U$. The vertices $w_{j}\in W$ are located at the leaf of the neighborhood trees and are selected by computing a path from the root node to the leaf node of the neighborhood trees. The computation of the path depends on the cumulative sum of degrees, $d_{u_{i},\omega}$, $\forall u_{i}\in U$ of each level of the neighborhood trees. Figure 1(b) shows one such path for selecting $w_{5}\in W$ for the right partition $K_{q}$ of $\delta$-clique $C_{q},q=1$. It is important to note that the selection of a vertex is based on the cumulative degrees $d_{u_{i},\omega}$ at every level of the neighborhood trees, i.e., $\sum_{u_{i}\in U}d_{u_{i},\omega}$, which could overlook a vertex $w_{j}\in W$ with high degree, i.e., with high number of neighbors, and thus, lowering the number of common neighbours found for the right partition of the $\delta$-clique. Additionally, to avoid the selection of the same vertex in the following iteration, the neighborhood trees need to be updated, which makes vertex selection a sequential process. Thus, to select vertices that lead to a larger set of common neighbors and to remove the dependency for the vertex selection, our algorithm selects vertices for the right partition of the $\delta$-cliques based on the non-increasing order of their degrees. We now describe the update required for the neighborhood trees to avoid repetitive selection of a vertex in $K_{q}$.
Updating neighborhood trees. With each selected vertex $w_{j}\in W$, the FM algorithm updates each neighborhood tree corresponding to all vertices $u_{i}\in U$ that have an edge with the selected vertex $w_{j}\in W$, i.e., neighborhood trees associated with $U_{K_{q}}=\{u_{i}\in U\ |\ w_{j}\in\Gamma(u_{i})\}$. This update is made by subtracting 1 from $d_{u_{i},\omega}$ at each node on the path from the root to the selected vertex $w_{j}$ (a leaf) regardless of whether the vertices $u_{i}\in U_{K_{q}}$ have an edge with $\forall w_{j}\in K_{q}$. Thus, we can say that with every clique the FM algorithm extracts, it eliminates the selection of at least one vertex $w_{j}\in W$ in future iterations of finding $\delta$-cliques. Thus restricting the extraction of cliques in future iterations even if there exists potential edges that can form non-trivial $\delta$-cliques (i.e., $\hat{k}>1$) improving compression. As we will explain in Section 5, this limitation is particularly relevant in less dense graphs, where there may not be enough common neighbors for the right partition of the $\delta$-clique to form additional cliques using the remaining edges. Therefore, in our approach we choose to update on the degree of each vertex $w_{j}$ in the right partition $K_{q}$ of the $\delta$-clique by subtracting the number of common neighbors found for $K_{q}$. This makes sure that only edges that will be removed by the $\delta$-clique are eliminated for the future iterations to find the $\delta$-clique, thereby giving the vertex $w_{j}\in W$ a chance to be selected for future $\delta$-cliques. We will now discuss the extraction of the $\delta$-cliques.
(a) FM
(b) CPGC
Figure 2: Progression of $\hat{m}$, $\hat{k}$, and number of cliques extracted for a graph with 128 vertices in each bi-partition, density 0.98, and $\delta=1$ by FM and CPGC.
Extracting $\delta$-cliques. A $\delta$-clique is formed by selecting $\hat{k}=k(n,\hat{m},\delta)$ vertices from $W$ in $G$. $\hat{k}$ depends on $\hat{m}$, the number of remaining edges, and on $n$, the number of vertices in $G$. It decreases as the number of iterations of the FM algorithm increases, as shown in Figure 2(a). With each iteration of the algorithm, $|U_{q}|\times|W_{q}|$ edges are removed from $G$ which results in decreasing $\hat{m}$, and thus $\hat{k}$. As more $\delta$-cliques are extracted, both $\hat{m}$ and $\hat{k}$ monotonically decrease. Furthermore, as we observe in Figure 2(a), $\hat{k}$ remains constant for several iterations. This is due to the fact that $\hat{k}$ is computed by using a function involving $\hat{m}$ and $n$, and the algorithm can remove a maximum of $n\times\hat{k}$ edges in each iteration. As a result, $\hat{k}$ is decreased by at most one, leading to a relatively slow decrease. Consequently, $\hat{k}$ remains constant for an increasing number of iterations as the rate of edges being removed decreases with decreasing $\hat{k}$, which motivates the selection of multiple $\delta$-cliques for the same $\hat{k}$ in each iterations of the algorithm. Figure 2(b), shows the progression of $\hat{m}$ and $\hat{k}$ and the number of $\delta$-cliques extracted by our algorithm, CPGC. We observe that for the same $\hat{k}$, unlike FM, CPGC extracts more $\delta$-cliques in one iteration leading to the formation of larger $\delta$-cliques compared to FM. For example, for $\hat{k}=4$, FM extracts one clique in each iteration i.e, from iteration 7 to 11, while CPGC extracts 8 cliques in one iteration. Extracting larger $\delta$-cliques results in removing more edges in each iteration of CPGC and thus accelerates the rate at which $\hat{m}$ decreases. Because of this we also observe that for the same graph CPGC extracted fewer $\delta$-cliques compared to FM. Despite of removing a smaller number of $\delta$-cliques, our experimental results (Section 5) show that CPGC obtains a better compression ratio than FM.
Our contributions. The above limitations of the FM algorithm (the best existing algorithm for clique partitioning-based graph compression) motivated us to design a novel algorithm called Clique Partitioning-based Graph Compression (CPGC). Our contributions are:
  1. 1.
    CPGC enables the removal of multiple $\delta$-cliques in each iteration, thereby obtaining a running time of $O(mn^{\delta})$ which is better than $O(mn^{\delta}\log^{2}n)$, the running time of FM.
  2. 2.
    CPGC enables the selection of vertices for forming cliques more than once. This can increase the number of cliques that are extracted during the compression, particularly in the case of large and high density graphs.
  3. 3.
    CPGC obtains an average compression ratio at least as good as that of the FM algorithm. The compression ratio is defined as $\frac{m}{m^{*}}$, where $m$ is the number of edges in the original graph $G$ and $m^{*}$ is the number of edges in the compressed graph $G^{*}$.
  4. 4.
    A comprehensive experimental analysis showing that on large dense graphs, CPGC achieves a compression ratio of up to 26% greater than that obtained by the FM algorithm and executes up to 105.18 times faster than the FM algorithm.
3 The Proposed Algorithm
3.1 Clique Partitioning-based Graph Compression (CPGC) Algorithm.
In this section, we present the design of our Clique Partitioning-based Graph Compression (CPGC) algorithm. For a given input graph $G$, CPGC builds the compressed graph $G^{*}$ of $G$ by partitioning $G$ into bipartite cliques. CPGC iteratively finds bipartite cliques in the given graph $G(U,W,E)$ and compresses it until finding new bipartite cliques does not contribute to the compression of $G$. The size of the right partition of the cliques is determined by $\hat{k}$, which guarantees a $\delta$-clique, and the selection of vertices $w_{j}\in W$ depends on the degree of vertices $d_{w_{j}}=|N(w_{j})|$. The partition of $G$ into bipartite cliques guarantees that each edge in $G$ is in exactly one clique, i.e., $E(C_{i})\cap E(C_{j})=\emptyset$, $\forall i,j$ and $i\neq j$, where $E(C_{i})$ is the set of edges of $\delta$-clique $C_{i}$. CPGC is given in Algorithm 1. The input of CPGC consists of the adjacency matrix $A$ of $G$, and a constant $\delta$.
Initialization (Line 1). For the given bipartite graph $G(U,W,E)$, CPGC initializes $q$, the index of the bipartite cliques extracted from $G$, the number of vertices, $n$ where $n=|U|=|W|$, the number of edges $\hat{m}$, and the degree of the vertices ${d_{w_{j}}}$ $\forall j=1,\ldots,|W|$, to $[0]_{n}$. CPGC also initializes the set of extracted bipartite cliques $\mathcal{C}$, to the empty set.
Algorithm 1 CPGC: Clique Partitioning-based Graph Compression Algorithm
1:$A$: Adjacency matrix;    $\delta$: a constant such that $0\leq\delta\leq 1$.
2:$q\leftarrow 0;\quad n\leftarrow|W|;\quad\hat{m}\leftarrow|E|;\quad{d_{w}}\leftarrow[0]_{n};\quad\mathcal{C}\leftarrow\emptyset$
3:for ${i=1,\ldots,n}$ do
4:    for ${j=1,\ldots,n}$ do
5:          ${d_{w_{j}}}\leftarrow{d_{w_{j}}}+E_{{u_{i}},w_{j}}$
6:$\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
7:while $\hat{k}>1$  do
8:    $\mathscr{C},\hat{A},\hat{d_{w}}\leftarrow\textsf{CSA(}q,\hat{k},A,n,d_{w}\textsf{)}$
9:    $\mathcal{C}\leftarrow\mathcal{C}\cup\mathscr{C}$
10:    $d_{w}\leftarrow\hat{d_{w}}$
11:    $A\leftarrow\hat{A}$
12:    $\hat{m}\leftarrow\sum_{j=1}^{|W|}d_{w_{j}}$
13:    $q\leftarrow q+|\mathscr{C}|$
14:    $\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
15:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$. This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
16:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
17:Output: $G^{*}$, the compressed graph of $G$.
@@ 3 The Proposed Algorithm
Partition size (Lines 2-5). CPGC computes the degree of each vertex $w_{j}\in W$, $d_{w_{j}}=|N(w_{j})|$, where $E_{{u_{i}},w_{j}}$ is 1, if there is an edge between ${u_{i}}$ and $w_{j}$, and 0, otherwise (Lines 2-4). It determines the size of the right partition of the bipartite clique, $\hat{k}$ (Line 5), which guarantees the existence of a $\delta$-clique with $k(n,\hat{m},\delta)$ [9]. It is important to say that $\hat{m}=|E|=\sum_{j=1}^{|W|}d_{w_{j}}$.
@@
Clique extraction (Lines 6-13). CPGC proceeds with extracting $\delta$-cliques until extracting new cliques does not contribute to the graph compression. This happens when $\hat{k}=1$, which results in obtaining trivial bipartite cliques. Thus, the while loop in Lines 6 to 13 is executed until trivial cliques are produced. Clique Stripping Algorithm (CSA), presented in Algorithm 2, finds $\delta$-cliques in the bipartite graph $G$. CSA takes the updated index of the bipartite clique, $q$, the size of the right partition of the bipartite clique, $\hat{k}$, the adjacency matrix, $A$, the size of the left partition of $G$, $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|,\ \forall w_{j}\in W$ as input. CSA, which is described in Subsection 3.2, returns the set of bipartite cliques $\mathscr{C}$ that it determined in the current execution, the updated adjacency matrix, $\hat{A}$, and the updated degree of vertices $\hat{d_{w}}$. In Line 8, CPGC adds the new bipartite cliques to the set of all bipartite cliques $\mathcal{C}$, updates the degree of vertices $d_{w}$ (Line 9), and adjacency matrix $A$ (Line 10). In Line 11, CPGC updates the number of edges, $\hat{m}$ in the given graph $G(U,W,E)$ after extracting the $\delta$-cliques, and the clique index $q$ (Line 12). Finally, in Line 13, after the $\delta$-cliques are removed, CPGC updates $\hat{k}$ for the next iterations.
@@
Graph compression (Lines 14-15). When the while loop (Lines 6-13) terminates, CPGC compresses the graph by adding a new vertex set $Z$ to the bipartite graph. Each vertex of the left and right partitions of clique $C_{q}$ are connected to a vertex $z_{q}$ via an edge, thus forming a tripartite graph. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The remaining edges in the given bipartite graph $G(U,W,E)$ connecting the set $U$ and $W$, which are not part of any $\delta$-clique are connected directly in the tripartite graph. Therefore, the compressed graph obtained by the CPGC preserves the path information of the original graph.
In Appendix A.2 we provide an example showing the execution of CPGC on a given bipartite graph.
3.2 Clique Stripping Algorithm (CSA).
@@
The Clique Stripping Algorithm (CSA)., given in Algorithm 2, extracts $\delta$-cliques from the given bipartite graph $G(U,W,E)$. CSA initially selects the vertices for the right partition $K_{c}$ of the clique $C_{c}$ and then the common neighbours for the set $K_{c}$. The input of CSA consists of the updated index of the clique, $q$, the size of the right partition $\hat{k}$ which guarantees a $\delta$-clique in $G$, the adjacency matrix $A$, the size of left/right partition $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|$.
@@
1:$q$: Index of bipartite clique;   $\hat{k}$: Size of the right bipartition in bipartite clique;
2:      $A$: Adjacency matrix;           $n$: Number of vertices in left/right bipartition of $G$;
3:      $d_{w}$: Degree of vertices in bipartition $W$.
4:$\mathcal{K}\leftarrow\emptyset$
5:Sort $d_{w}$ in non-increasing order.
6:Let $d_{w_{\pi(1)}},d_{w_{\pi(2)}},\ldots,d_{w_{\pi(n)}}$ be the order.
7:$j\leftarrow 1$
8:while $d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}$ do
9:    $\mathcal{K}\leftarrow\mathcal{K}\cup\{w_{\pi(j)}\}$
10:    $j\leftarrow j+1$
11:$\gamma\leftarrow\big{\lfloor}{\frac{|\mathcal{K}|}{\hat{k}}}\big{\rfloor}$
12:for  $c=q+1,\ldots,q+\gamma$ do
13:    $K_{c}\leftarrow\{w_{\pi_{(j)}}\in\mathcal{K}\ |\ (c-(q+1))\cdot\hat{k}<j\leq(c-q)\cdot\hat{k}\}$
14:    $U_{K_{c}}\leftarrow\{u_{i}\in U|K_{c}\subseteq N(u_{i})\}$
15:    $C_{c}\leftarrow\{(U_{K_{c}},K_{c})\}$
16:    Update $A$, by removing edges $(U_{K_{c}}\times K_{c})\in C_{c}$
17:    Update $d_{w}$ by subtracting $|U_{K_{c}}|$ from each $d_{w_{\pi(j)}}$, where $w_{\pi(j)}\in K_{c}$
18:$\mathscr{C}\leftarrow\{C_{q+1},\ldots,C_{q+\gamma}\}$
19:Output Set of cliques $\mathscr{C},\hat{A},\hat{d_{w}}$
@@
Initialization (Lines 1-3). CSA initializes the set of selected vertices for clique extraction, $\mathcal{K}$ to the empty set. In Line 2, it sorts $d_{w}$ in non-increasing order of the degrees of vertices $w_{j}\in W$, i.e., $d_{w_{\pi(1)}}\geq d_{w_{\pi(2)}}\geq\ldots\geq d_{w_{\pi(n)}}$. Then, in Line 3 it initializes the index $j$ in the list of sorted vertices to 1.
Vertex selection (Lines 4-7). CSA selects the vertices of a clique to be extracted, based on their degrees. Selecting a vertex with higher degree results in a larger set of neighbours that can be part of the selected clique. Therefore, CSA selects the vertices whose degrees are greater than or equal to the $\hat{k}$-th largest degree in the sorted $d_{w}$, i.e., $\mathcal{K}=\{w_{\pi(j)}\ |\ d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}\}$. CSA adds the selected vertices to $\mathcal{K}$ and increments the iterator $j$ (Lines 5-6). Thus, the while loop in Lines 4- 6 is executed until all vertices with degrees greater than $d_{w_{\pi(\hat{k})}}$ are chosen. In Line 7, CSA calculates $\gamma$ to partition the set $\mathcal{K}$ such that the size of a partition is less than or equal to $\hat{k}$. Therefore, every time CSA is executed it finds $\gamma$ cliques.
@@
Clique extraction (Lines 8-13). The for loop (Lines 8-9) first partitions $\mathcal{K}$ into subsets $K_{c}$, such that $\bigcup_{c=q+1,\ldots,q+\gamma}K_{c}=\mathcal{K}$ and $K_{q+i}\bigcap K_{q+j}=\emptyset,\ i=1,\ldots,\gamma,j=1,\ldots,\gamma,i\neq j.$ Then it constructs $U_{K_{c}}$ such that each vertex in $U_{K_{c}}$ is part of the set of common neighbors of $K_{c}$ (Line 10). In Line 11, CSA forms clique $C_{c}$ with partitions $U_{K_{c}}$ and $K_{c}$, where $|U_{K_{c}}|\geq\lceil{n^{1-\delta}\rceil}$ and $|K_{c}|=\hat{k}$. After forming the clique $C_{c}$, in Line 12, the algorithm updates the adjacency matrix by removing the edges in the clique $C_{c}$ and in Line 13, updates the degrees of the vertices that are part of the clique by removing the size of left partition from the degree of each vertex $w_{\pi(j)}\in K_{c}$, i.e., $d_{w_{\pi(j)}}=d_{w_{\pi(j)}}-|U_{K_{c}}|,\ \forall\ w_{\pi(j)}\in K_{c},\ c=q+1,\ldots q+\gamma$. Finally, in Line 14, CSA forms $\mathscr{C}$, the set of all the cliques extracted in the current execution, and in Line 15, it returns the set of cliques $\mathscr{C}$, the updated adjacency matrix $\hat{A}$, and the updated degrees of vertices $\hat{d}_{w}$ to CPGC.
Time complexity of CSA. CSA takes $O(n\log n)$ to sort the degrees of vertices $d_{w}$, in Line 2. The while loop in Lines 4-6, takes at most $O(n)$ to select vertices for the $\delta$-cliques. The for loop (Lines 8-13) executes $O(\gamma)$ times, where $1\leq\gamma\leq\frac{n}{\hat{k}}$. In the for loop, Line 9 and 10 takes $O(\hat{k})$ and $O(n\hat{k})$ to find the right and left partition of the $c$-th $\delta$-clique, respectively. CSA takes $O(n\hat{k})$ to remove the edges in the $c$-th $\delta$-clique, in Line 12. It takes $O(\hat{k})$ to update the degrees $d_{w}\in K_{c}$, in Line 13. Therefore, the total running time of CSA is dominated by Lines 10 and 12 and thus CSA takes $O(n\gamma\hat{k})$ time.
Time complexity of CPGC. CPGC takes $O(n^{2})$ to calculate the degrees $d_{w}$ of vertices in $W$ in Lines 2-4. The running time of the while loop (Lines 6-13) is dominated by the running time of the function CSA in Line 7 which is given by Algorithm 2. CSA takes $O(n\gamma\hat{k})$ to remove $\gamma\hat{k}n^{1-\delta}$ edges. Therefore, on average it takes $O(n^{\delta})$ time to remove one edge from the given graph. Thus, the while loop in CPGC takes $O(mn^{\delta})$ time to extract $\delta$-cliques. In the end, to compress the graph, CPGC in Lines 14-15, takes linear time. Therefore, the total running time of CPGC is $O(mn^{\delta})$ which is dominated by the while loop in Lines 6-13.
4 Properties of CPGC
Theorem 4.1

### v3

@@ 3.1 Clique Partitioning-based Graph Restructuring (CPGR) Algorithm
[…]
8:for $i=1,\ldots,n$ do
9:  for $j=1,\ldots,n$ do
10:    ${d_{w_{j}}}\leftarrow{d_{w_{j}}}+E_{{u_{i}},w_{j}}$
[…]
12:while $\hat{k}>1$ do
13:  ($\mathscr{C},\hat{A},\hat{d_{w}})\leftarrow\textsf{CSA(}q,\hat{k},A,n,d_{w}\textsf{)}$
14:  $\mathcal{C}\leftarrow\mathcal{C}\cup\mathscr{C}$
15:  $d_{w}\leftarrow\hat{d_{w}}$
16:  $A\leftarrow\hat{A}$
17:  $\hat{m}\leftarrow\sum_{j=1}^{|W|}d_{w_{j}}$
18:  $q\leftarrow q+|\mathscr{C}|$
19:  $\hat{k}\leftarrow\Big\lfloor{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big\rfloor$
[…]
21:for $i\in\mathcal{C}=\{\mathscr{C}_{1}(U_{1},W_{1}),\ldots,\mathscr{C}_{q}(U_{q},W_{q})\}$ do
22:  $Z\leftarrow Z\cup\{z_{i}\}$
23:  $E^{*}\leftarrow E^{*}\cup\{(u,z_{i}),(z_{i},w)\}\quad\forall u\in U_{i},\quad\forall w\in W_{i}$
24:  $\mathcal{E}_{\mathcal{C}}\leftarrow\mathcal{E}_{\mathcal{C}}\cup E_{i}$
[…]
26:Output: $G^{*}(U,Z,W,E^{*})$, the restructured graph of $G$.
@@
Partition size (Lines 6-9). CPGR computes the degree of each vertex $w_{j}\in W$, $d_{w_{j}}=|N(w_{j})|$, where $E_{{u_{i}},w_{j}}$ is 1, if there is an edge between $u_{i}$} and $w_{j}$, and 0, otherwise (Lines 6-8). It determines the size of the right partition of the bipartite clique, $\hat{k}$ (Line 9), which guarantees the existence of a $\delta$-clique with $k(n,\hat{m},\delta)$ [14]. It is important to say that $\hat{m}=|E|=\sum_{j=1}^{|W|}d_{w_{j}}$.
Clique extraction (Lines 10-17). CPGR proceeds with extracting $\delta$-cliques until extracting new cliques does not contribute to the reduction in the number of edges. This happens when $\hat{k}=1$, which results in obtaining trivial bipartite cliques. Thus, the while loop in Lines 10 to 17 is executed until trivial cliques are produced.
@@
Clique Stripping Algorithm (CSA), presented in Algorithm 2, finds $\delta$-cliques in the bipartite graph $G$. CSA takes the updated index $q$ of the bipartite clique, the size $\hat{k}$ of the right partition of the bipartite clique, the adjacency matrix $A$, the size $n$ of the left partition of $G$, and the degree $d_{w_{j}}=|N(w_{j})|,\ \forall w_{j}\in W$ as input. CSA, which is described in Subsection 3.2, returns the set $\mathscr{C}$ of bipartite cliques that it determined in the current execution, the updated adjacency matrix $\hat{A}$, and the updated degree of vertices $\hat{d_{w}}$. n Line 12, CPGR adds the new bipartite cliques to the set of all bipartite cliques $\mathcal{C}$, updates the degree of vertices $d_{w}$ (Line 13), and adjacency matrix $A$ (Line 14). In Line 15, CPGR updates the number of edges, $\hat{m}$ in the given graph $G(U,W,E)$ after extracting the $\delta$-cliques, and the clique index $q$ (Line 16). Finally, in Line 17, after the $\delta$-cliques are removed, CPGR updates $\hat{k}$ for the next iterations.
@@
Graph restructuring (Lines 18-23). When the while loop (Lines 10-17) terminates, CPGR restructures the graph by adding a new vertex set $Z$ to the bipartite graph. Each vertex in the left and right partitions of clique $\mathscr{C}_{q}$ is connected to a new vertex $z_{q}$, reducing the number of edges from $|U_{q}|\times|W_{q}|$ to $|U_{q}|+|W_{q}|$ (Lines 18-23). In Line 22, $\mathcal{E}_{\mathcal{C}}$ is updated with edges from the extracted $\delta$-cliques. Remaining edges in the bipartite graph $G(U,W,E)$, not part of any bipartite $\delta$-cliques, are connected directly in the tripartite graph (Line 23). In Line 24 the algorithm returns the restructured graph $G^{*}(U,Z,W,E^{*})$ having a reduced number of edges.
In Appendix A.2 we provide an example showing the execution of CPGR.
3.2 Clique Stripping Algorithm (CSA)
@@
The Clique Stripping Algorithm (CSA), given in Algorithm 2, extracts $\delta$-cliques from the given bipartite graph $G(U,W,E)$. CSA initially selects the vertices for the right partition $K_{c}$ of the clique $C_{c}$ and then the common neighbors for the set $K_{c}$. The input of CSA consists of the updated index $q$ of the clique, the size $\hat{k}$ of the right partition which guarantees a $\delta$-clique in $G$, the adjacency matrix $A$, the number of vertices $n$, and the degree $d_{w_{j}}=|N(w_{j})|$ .
@@
1:$q$: Index of bipartite clique;
2:  $\hat{k}$: Size of the right bipartition in bipartite clique;
3:  AA: Adjacency matrix;
4:  $n$: Number of vertices in left/right bipartition of $G$;
5:  $d_{w}$: Degree of vertices in bipartition $W$.
[…]
7:Sort $d_{w}$ in non-increasing order. Let $d_{w_{\pi(1)}},d_{w_{\pi(2)}},\ldots,d_{w_{\pi(n)}}$ be the order.
[…]
9:while $d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}$ do
10:  $\mathcal{K}\leftarrow\mathcal{K}\cup\{w_{\pi(j)}\}$
11:  $j\leftarrow j+1$
[…]
13:for $c=q+1,\ldots,q+\gamma$ do
14:  $K_{c}\leftarrow\{w_{\pi_{(j)}}\in\mathcal{K}\ |\ (c-(q+1))\cdot\hat{k}<j\leq(c-q)\cdot\hat{k}\}$
15:  $U_{K_{c}}\leftarrow\{u_{i}\in U|K_{c}\subseteq N(u_{i})\}$
16:  $C_{c}\leftarrow\{(U_{K_{c}},K_{c})\}$
17:  Update AA, by removing edges $(U_{K_{c}}\times K_{c})\in C_{c}$
18:  Update $d_{w}$ by subtracting $|U_{K_{c}}|$ from each $d_{w_{\pi(j)}}$, where $w_{\pi(j)}\in K_{c}$
[…]
20:Output ($\mathscr{C},\hat{A},\hat{d_{w}}$)
@@
Initialization (Lines 1-3). CSA initializes the set $\mathcal{K}$ of selected vertices for clique extraction to the empty set. In Line 2, it sorts $d_{w}$ in non-increasing order of the degrees of vertices $w_{j}\in W$, i.e., $d_{w_{\pi(1)}}\geq d_{w_{\pi(2)}}\geq\ldots\geq d_{w_{\pi(n)}}$. In Line 3 it initializes the sorted vertices list index $j$ to 1.
Vertex selection (Lines 4–7). CSA selects vertices for clique extraction based on their degrees, since vertices with higher degree are more likely to share a larger set of common neighbors and thus participate in dense bipartite cliques. Rather than selecting exactly $\hat{k}$ vertices, CSA constructs a candidate set $\mathcal{K}=\{w_{\pi(j)}\mid d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}\}$, which contains all vertices whose degree is at least the $\hat{k}-$th largest degree in the sorted sequence $d_{w}$. As a result, $|\mathcal{K}|\geq\hat{k}$ in general. CSA adds these vertices to $\mathcal{K}$ by iterating through the sorted order and incrementing the index $j$ (Lines 5–6). Thus, the while-loop in Lines 4–6 ensures that all vertices tied at the $\hat{k}-$th degree threshold are included, avoiding arbitrary tie-breaking and forming a maximal candidate set for the right partition.
This threshold-based selection increases the likelihood of identifying a larger set of common neighbors in the left partition, thereby improving the quality of the extracted $\delta$-cliques. In Line 7, CSA computes $\gamma=\lfloor|\mathcal{K}|/\hat{k}\rfloor$, which determines how many $\delta$-cliques can be extracted. The set $\mathcal{K}$ is then partitioned into $\gamma$ subsets of size $\hat{k}$ (Line 9), allowing CSA to extract multiple $\delta$-cliques in a single execution.
@@
Clique extraction (Lines 8-13). The for loop (Lines 8-9) first partitions $\mathcal{K}$ into subsets $K_{c}$, such that $\bigcup_{c=q+1,\ldots,q+\gamma}K_{c}=\mathcal{K}$ and $K_{q+i}\bigcap K_{q+j}=\emptyset,\ i=1,\ldots,\gamma,j=1,\ldots,\gamma,i\neq j.$ Then it constructs $U_{K_{c}}$ such that each vertex in $U_{K_{c}}$ is part of the set of common neighbors of $K_{c}$ (Line 10). In Line 11, CSA forms clique $C_{c}$ with partitions $U_{K_{c}}$ and $K_{c}$, where $|U_{K_{c}}|\geq\lceil{n^{1-\delta}\rceil}$ and $|K_{c}|=\hat{k}$. After forming the clique $C_{c}$, in Line 12, the algorithm updates the adjacency matrix by removing the edges in the clique $C_{c}$ and in Line 13, updates the degrees of the vertices that are part of the clique by removing the size of left partition from the degree of each vertex $w_{\pi(j)}\in K_{c}$, i.e., $d_{w_{\pi(j)}}=d_{w_{\pi(j)}}-|U_{K_{c}}|,\ \forall\ w_{\pi(j)}\in K_{c},\ c=q+1,\ldots q+\gamma$. Finally, in Line 14, CSA forms $\mathscr{C}$, the set of all the cliques extracted in the current execution, and in Line 15, it returns the set of cliques $\mathscr{C}$, the updated adjacency matrix $\hat{A}$, and the updated degrees of vertices $\hat{d}_{w}$ to CPGR.
[…]
4 Properties of CPGR
The transformation performed by CPGR preserves vertex-to-vertex reachability and the connected components of the original graph. Consequently, the existence of paths between any pair of vertices in $G$ is maintained in the restructured graph $G^{*}$. However, CPGR does not necessarily preserve local structural properties such as vertex degrees, neighborhood structure, or exact shortest-path distances, since edges belonging to $\delta$-cliques are replaced by paths of length two. We formalize this preservation of path information in the following theorem.
Theorem 1.